What is the probability of getting a perfect square from the numbers 1 to 100?




A.


1/4



B.


1/10



C.


1/25



D.


1/5



E.


None of the above



The probability that two events, A and B, will both occur is represented by the following multiplication rule:




A.


P(A ∩ B) = P(A│B) × P(B)



B.


P (A ∩ B) = P(A) × P(B)



C.


P(A ∩ B) = P(B│A) × P(B)



D.


Both B and C.



E.


None of the above




1
Expert's answer
2022-05-29T16:10:53-0400

1. Perfect squares: "1,4,9,16,25,36,49,64,81,100."

The probability of getting a perfect square from the numbers 1 to 100 is


"\dfrac{10}{100}=\dfrac{1}{10}"

B. 1/10


2.

A.


P(A ∩ B) = P(A│B) × P(B)